OpenStudy (anonymous):

i mean what is 0.54 divided by 0.756

OpenStudy (anonymous):

Yes, it's 0.7142857143 which is 5/7.

OpenStudy (anonymous):

Since for some reason you don't believe me, here's the long-winded version: \[\Large \begin{array}{l} \frac{{0.54}}{{0.756}} = 0.7142857143\\\\ 0.54 = \frac{{27}}{{50}}\\ 0.756 = \frac{{189}}{{250}}\\\\ \frac{{27}}{{50}} \div \frac{{189}}{{250}} = \frac{{27 \times 250}}{{50 \times 189}} = \frac{{6750}}{{9450}}\\\\ \frac{{6750}}{5} = \frac{{9450}}{7}\\\\ \frac{{6750}}{{9450}} = \frac{5}{7} \end{array}\]

OpenStudy (anonymous):

\[\frac{54}{100}/\frac{756}{1000}=\frac{5}{7} \]